As of September 2021, the population (16+ years of age) in the U.S. is 263

million people. The labor force participation rate is 61%, and the unemployment rate

is 4.8%. What is the employment level?


1
Expert's answer
2022-03-25T15:08:00-0400

Employment rate = Divide the number of employed people by the total labor force multiply by 100.

Labour Participation= labour force÷ Total working age population

Let labour force be X

61%= "\frac{X}{263}\times 100"

X= "\frac{263\times 61}{100}= 160.43"

Let unemployed be Y

Number of unemployed people= "\frac{160.43\times4.8}{100}= 7.70064"

Employed people= 160.43- 7.70064= 152.72936